OpenStudy (snowflakelove15):

You can approximate e by substituting large values for n into the expression ________. A. (1-n)^1/n B.(1+1/n)^n C. (1+n)^1/n D.(1-1/n)^n

OpenStudy (zzr0ck3r):

google

OpenStudy (evoker):

I would recommend looking at the wiki for e

OpenStudy (snowflakelove15):

I did I couldn't find anything

OpenStudy (evoker):

It's in the first couple sentences on the Wikipedia entry

OpenStudy (zzr0ck3r):

B

OpenStudy (zzr0ck3r):

lol

jimthompson5910 (jim_thompson5910):

You can plug in large numbers into n (like n = 1000) and see which get you close to e = 2.71828182846...

OpenStudy (snowflakelove15):

Thank you both of you, B was correct!

OpenStudy (zzr0ck3r):

hint: \(e^x=\lim_{n\rightarrow \infty}(1+\frac{x}{n})^n\)

jimthompson5910 (jim_thompson5910):

Yes it's B since (1+1/n)^n = (1+1/1000)^1000 = 2.7169239322356 which is fairly close. Just off at the 3rd decimal digit. Use larger values of n to get closer

OpenStudy (zzr0ck3r):

https://en.wikipedia.org/wiki/List_of_representations_of_e This has a bunch of them